Problem 6: On a very hot summer day, 5 percent of the production employees at Midland Stalies Steel are abit from work. The production employees are to be selected at random for a special in-depth study on absenteeism a) What is the probability of selecting 10 production none of them are absent? employees at random on a hot summer day and finding tha What is the probability of selecting 10 production employees at random on a hot summer day and finding at least one was absent? b) Problem 7. A shipment of 12 microwave ovens contains three defective units. A vending company has ordered four of these units, and because each is identically packaged, the selection will be random. What is the probability that (a) all four units are good, (b) exactly two units are good, and (c) at least two units are good?
Answer #1

Solution:-
6)
a) B(10,0,0.05) = 0.95^10 = 0.5987

b) => 1-0.5987 = 0.4013

7)
a. 9C4/12C4 = 126/495 = 0.2545

b. 9C2*3C2/12C4 = 108/495 = 0.2182

c. 486/495 = 0.9818